22. Much of the demand for U.S. agricultural output has come from other countries. In 1998,

the total demand for wheat was Q = 3244 – 283P. Of this, total domestic demand was QD =

1700 – 107P, and domestic supply was QS = 1944 + 207P. Suppose the export demand for

wheat falls by 40 percent.

a. U.S. farmers are concerned about this drop in export demand. What happens to the

free-market price of wheat in the United States? Do the farmers have much reason to worry?

b. Now suppose the U.S. government wants to buy enough wheat to raise the price to

$3.50 per bushel. With the drop in export demand, how much wheat would the

government have to buy? How much would this cost the government?


1
Expert's answer
2021-02-22T14:02:42-0500

1)Firstly we have to find initial export demand. Qd(ex)=3244-283P-(1700-107P)=1544-176P. After 40% fall it becomes 0.6×(1544-176P)=926.4-105.6P. Equilibrium price before demand fall can be found from equation 3244-283P=1944+207P. 1300=490P; P=2.65. Equilibrium price after fall is 1700-107P+926.4-105.6P=1944+207P; 682.4=419.6P; P=1.63. So, equilibrium price reduced by $1.02 till $1.63, thus farmers have reason to worry.

2. Suppose X is the quantity that government should buy. Then total demand is 1700-107P+926.4-105.6P+X. At the price of 3.50 it equals to 2626.4- 744.1+X =1882.3+X. Total supply at the price of $3.50 is 1944+207×3.5=2668.5. Then, 1882.3+X=2668.5; X=786.2. It will cost for government 786.2×3.5=2751.7